Jessica is driving on the highway. She begins the trip with 14 gallons of gas in her car. The car uses up one gallon of gas every30 miles.
Let g represent the number of gallons of gas she has left in her tank, and letd represent the total distance (in miles) she has traveled

Answer by ankor@dixie-net.com(22740)

begins the trip with 14 gallons of gas in her car.
The car uses up one gallon of gas every 30 miles.
Let g represent the number of gallons of gas she has left in her tank, and let d
represent the total distance (in miles)
:
d = 30(14 - g)
:
For example, if she uses two gallons, there will be 12 gal left in the tank; g=12
d = 30(14 - 12)
d = 30(2)
d = 60 mi